# Reviewing Intelligent Cinematography: AI research for camera-based video production

### v1

@@ 2.5 Automated 3-D Capture
Automated 3-D capture is a popular topic in IC research, whether for analysis or for solving the inverse graphics problem. Here, the objective is to discover a 3-D representation using 2-D imagery and/or other sensor data. In Figure 4, we illustrate simplified representations of the current approaches to the problem.
@@
(b) Neural Radiance Fields sample points in space to estimate a colour and density value for each sample projected along a pixel-ray
(c) Gaussian Splatting uses a point cloud to estimate the position of samples in space and uses this to estimate the colour, density scale and rotation of each Gaussian point
@@
A popular capture method is Laser Informed Photogrammetry (e.g. LiDAR) which creates 3-D point-clouds from measurements of point depth relative to a camera’s position. This can now be done on a mobile phone thanks to the introduction of mobile LiDAR camera sensors. These methods are promising for capturing enclosed spaces, however there are challenges with processing information over greater distances as well as limitations associated with noisy data. Additionally, these cases require photogammetric sensors which limits applicability when only image/video data is available.
@@
Recently, neural radiance field (NeRF) methods have been used to solve this problem, whereby the aim is to model the rendering processes involved with automated capture. Unlike Photogammtery which may represent scenes as 3-D structures (mesh, voxel-grid, point cloud), neural methods represent scene properties as neural representations which are sampled to interpolate views defined by the camera extrinsics in a virtual space. Given a set of images describing a real environment, NeRF networks are expected to reliably learn the visual features of a scene.
@@
NeRFs represent an important breakthrough for cinematographers, as can be adapted to a range of cinematographic tasks. For example, rendering 3-D from 2-D images can avoid costly re-shoots caused by poor lighting, scenery, weather and acting deficiencies [167, 255, 287]. Shots can also be re-worked in post production, meaning that shot-type, camera jitter, pacing and focus can also be modified. While the field of NeRF research is relatively new, it has developed rapidly and can now deliver high quality, compact solutions, leading us to expect production-ready NeRFs in the near future.
The NeRF model samples volume density, $\sigma$, and colour radiance, $\mathbf{c}$, provided a 5-D input comprised of 3-D co-ordinates, $\mathbf{o}\;\epsilon\;\mathbb{R}^{3}$, plus 2-D viewing direction, $\mathbf{d}\;\epsilon\;\mathbb{R}^{2}$, which represents the position and viewing direction of a sample in space. A sample can be thought of as either a volumetric line-segment along a ray or a voxel intersected by a ray casted into the scene. Simply put, for each pixel $(x,y)$ in an image a ray vector $\mathbf{r_{x,y}}=\mathbf{o}+t\mathbf{d_{x,y}}$ exists, where $t=0$ represents the focal point of an image, $t=n$ represents the position of the image plane along the ray and $t>n$ represents point samples along a ray where $n$ sets the scalar distance from the focal point to the image plane and can vary w.r.t lens distortion. We refer the reader to the Nerfstudio documentation [243] [^1] Accessible online: https://docs.nerf.studio/en/latest/nerfology/model_components/visualize_cameras.html which overviews the different types of camera models, sampling schemes and sample representations that are found in the NeRF literature.
[…]
$\hat{\mathbf{C}}(\mathbf{r})=\sum^{t_{far}}_{i={t_{near}}}(1-\exp(-\sigma_{i}\delta_{i}))\mathbf{c_{i}}\exp({\sum^{i-1}_{j=t_{near}}-\sigma_{j}\delta_{j}})$ (2)
Subsequently, a loss function $\mathcal{L}(\mathbf{C^{*}},\hat{\mathbf{C}})$ is used to optimise the predicted ray colour w.r.t to the colour of a ground truth pixel, $\mathbf{C^{*}}$.
To reduce the influence from spectral bias [202] NeRF maps position and viewing directions to a view-point’s fourier features, using the encoding $\gamma$ in Equation 3, where $k$ is a hyper-parameter defining the dimensionality of the feature vector (the bandwidth of our discretized frequency encoding). It is well known in neural representation research that coordinate-MLPs struggle to learn high-frequency signal details, hence the need for encoding frequencies using $\gamma$. However there are several studies that discuss alternative changes to the MLP activation functions which forgo the need for discretized frequency encoding [233, 224]. For example, Saragadam et al [224] proposes an MLP using wavelet activation (called the wavelet implicit representation representation (WIRE)) while Sitzmann et al [233] proposes a sinusoidal activation (called sinusoidal implicit representation (SIREN)). These are shown to not only reduce the size of the MLP but also capture a higher-bandwidth of frequencies. However, despite Sitzmann et al [233] and Saragadam et al [224] showcasing higher quality and faster convergence, they discuss how un-actuated these results are in NeRF research.
@@
Several approaches have been proposed to speed up NeRF computation: multi-resolution hash encodings, voxel grids and voxel-trees [171, 262, 283, 133]. This is particularly relevant for researchers interested in the dynamic NeRF paradigm, as model complexity increases significantly when modelling 4-D space and time. The PlenOctrees representation [283], illustrated in Figure 5, builds off NeRF’s ability to functionally represent a view-dependant scene by, (1) representing the structure of a scene as a sparse-voxel octree [268, 220, 133], and (2) sampling colour by involving spherical harmonics. (1) facilitates the rendering step through fast-to-access representations (octrees). Whereas (2) allows us to model view-dependency as octree leaves by mapping the view dependant colour values to a sphere’s surface and predicting the coefficients of the spherical harmonic equation (with fixed degree) to return a colour value where the pixel-ray intersects the sphere. Subsequently, the rendering speed and learning is optimised significantly, and one could argue novel-view quality has improved too. One caveat is the amount of data a representation will consume (in the magnitude of gigabytes), which perhaps is acceptable for those with required hardware, however there exist more affordable explicit options [45, 212]. Additionally, there is some difficulty with visualising unbounded scenes as voxel-grids occupy finite space. Yu et al [283] overcomes this by using NeRF++ [290] for rendering out-of-bound scenery. This is accomplished by modelling the foreground and background as separate components.
Figure 5: Method for neural rendering of 3-D images presented in Yu et al [283]
@@
With similar motivations, 3-D Gaussian Splatting (GS) has been proposed by Kerbl et al [126] as a means of significantly reducing rendering time by using: (1) a point cloud representation of Gaussian “blobs” with position, covariance, colour and opacity properties, reducing the unnecessary computation of empty space present in NeRF models and (2) tile splatting for rendering. The covariance property is represented as a $3\times 3$ matrix $\Sigma$ in Equation 4 and determines the scale $S$ and rotation $R$ of the blob in space. In practice, $\Sigma$ is approximated to a $2\times 2$ matrix, avoiding the last row and column in the original matrix, leading to faster computation.
$\Sigma=RSS^{T}R^{T}$ (4)
@@
Additionally tile splatting also speeds up rendering with the following steps. We highlight the importance of the sorting approach, (3), as this limits NeRFs from achieving similar computational goals.
@@
Finally, to train a GS model a scene is initialised with a sparse set of point-clouds that are “densified” during training by cloning, re-scaling and re-positioning blobs to fit the geometry corresponding to a set of training view/image. Blobs that are essentially transparent (low $\alpha$ contribution) are pruned.
Conclusively, GS maybe selected over NeRFs due to their enhanced computational abilities, and in many cases improved performance. However for the dynamic paradigm, the current set of benchmark scenes only last up to a minute. As existing dynamic GS models are much less compact than NeRF alternatives, this presents a challenge for storing and sharing 4-D models. Additionally, as GS is a more recent development than NeRF we are yet to see the same attention towards cinematographic based tools, such as scene and camera editing or representational transformations such as GS to mesh. Though, the research space is adapting at a fast pace so we believe it is only a matter of time before we see these limitations addressed.
Finally, we underpin the importance of social awareness and responsibility involved with this field. Unlike generative AI, we notice almost no attention to the ethical problems that may arise from creating digital doubles of real humans. We urge researchers involved with this field to take more of an active role raising awareness.
@@ 3.1 General Production Applications
In this section we present research relating to camera management and control, visual analysis, cinematographic assistance and workflow optimization.
@@
3.1.1 Computational Language Structures for Scene Analysis, Automated Labelling Schemes and Camera Management
@@
Cinematographic language is used to communicate the state of production and relevant processes. It enables production staff to effectively communicate and react coherently to unforeseen dilemma. IC applications require a similar level of attention. Cinematic theory is often formalized for tasks such as automated camera control and scene analysis. For example, in Jhala and Young [115] state machines are used to model cinematographic shots for camera planning in a virtual environment. In this context, in this section we discuss current approaches to human-AI interaction, with particular reference to automated camera control; we refer to this as idiomatic capture.
@@
A heuristic approach to camera control was introduced by Jhala and Young [114], based on the decompositional partial order causal link planning algorithm [280]. This formalizes idiomatic capture by linking an approximated scene representation to a predetermined set of camera control responses. A conflict resolution algorithm ranks known actions with conditional operators to determine the optimal action and respective duration. To express idioms, Jhala and Young [114] distinguishes a set of four requirements to mediate contextual differences between different productions: (1) Story Representation, (2) Real World Representation, (3) Rhetorical Coherence and (4) Temporal Consistency. The first two points describe the physical and contextual landscapes of a set, hence it is necessary to have both geometric (or visual) and semantic representations of a production. The third point expresses the need for rhetorical structure to ensure actions are executed decisively, e.g. a hierarchical-model for selecting shot-types. The final point specifies consistency with regards to the temporal aspect of filming.
@@
Similarly, the Declarative Camera Control Language (DCCL) uses a heuristic-based decision tree for idiomatic capture [55], Figure 6. This method sets forth a hierarchical structure for automated shot-composition and camera control by breaking down scenes into idiom-specific frame sequences. Here, the relationship of consecutive frame sequences is dependant on temporal links between idioms. A heuristic evaluator then selects a candidate action by scoring possible responses and evaluating decisions via a decision tree.
@@
Figure 6: A three stage pipeline which uses idioms defined by camera and actor states as inputs to a DCCL compiler. Selected frames are evaluated w.r.t to the “film tree”
@@
Idiomatic language structures have also been used for shot labelling [215] and evaluation [285]. In more recent times, the Prose Storyboard Language (PSL) [215] has been proposed as a high-level language for describing shots relative to visual composition and idiomatic camera control. This is demonstrated with Alfred Hitchcock’s North by Northwest, where a prompt not only describes the shot-type and fore, mid and background compositions (separately) but the type of transition as well. As PSL uses AND/OR decision trees it can be easily configured to add/modify labels. While the authors have not tested this on an automated labelling scheme, there is cause for such investigation as idiom-based languages might overlook newly adopted idioms and foreign capture practices.
@@
Embedded Constrained Patterns (ECP) is a dedicated query language, conscious of physical cinematographic limitations [272]. Unlike PSL, ECP is more comprehensive in its description - labelling montages rather than individual shots. Labels are assigned using the following descriptors:
@@
Wu and Christie [272] propose ECP alongside a automated search algorithm for optimal shot-sequence description. The search method is partitioned into two stages, (1) build a cache of valid solutions, and (2) apply a double recursive depth search to choose optimal description from the set of valid solutions. During stage (1) valid solutions are separated into three sets: (i) FC - a set of frames which satisfies framing constraints defined by ECP, (ii) RC - a set of frame couples, $[f_{i},f_{i+1}]$ which satisfies the ECP’s relational constraint, and (iii) SC - a set of frame sequences, $s_{m}=[f_{l},f_{F}]$, where $f_{l}$ and $f_{F}$ represent the start and end frames of the frame sequence. In the second stage, the search process iterates over all frames in each sequence, $f_{i}\epsilon s$, where subsequent frames are validated as part of the local sequence, $s$, or global sequence, $S$.
Alongside the technical implementation of ECP, the authors have plans for integration with Unity 3-D to apply ECP to their montages; which we see as a positive yet infrequent consideration in IC. It would be interesting to reverse the order of use. Whereby, rather than providing linguistic descriptors of a 3-D environment we could trigger camera actions through prompts written in PSL or ECP.
Additionally, we find that languages built for human-AI interaction are not easily comparable. The choice of language structure can vary drastically depending on the task for automation. While there is no current standard for general cinematography, we observe interest from Movie Labs, who look to define a set of computational linguistic structures as a standard for cloud-based workflow optimization [132]. The shift to cloud-based computing is indicative of another paradigm of production. Notably, as assets and tools are usually stored or executed offline changes to shared resources can be difficult to monitor on one platform. This could be facilitated by file management approaches which mimic, for example, code-sharing practices. However, this may be difficult to scale for productions requiring a large number of assets or productions which involve external international collaborators who may be un-knowledgeable of standard practices.
@@ 3.1.2 Directive Assistants
Converting most cinematographic concepts into controllers presents a challenging task [56]. Notions of shot compositions, shot types and shot transitions are bound by real world problems such as the cost of production, physics and the topology of a set. Thus, AI directive assistants (DAs) have been introduced to alleviate these production challenges. Current IC research addresses tasks such as deriving shot lists, shot plans, optimising camera placement and controlling robotic camera rigs.
@@
There is a tendency to using semantic representations of a work environment to provoke DAs. For example, through DCCL, Christianson et al [55] demonstrates a heuristic approach that relies on idiomatic-based practice for decision making on shot composition. Similarly, He et al [94] discuss a different heuristic approach, utilizing a set of finite state machines [^2] A state-based control architecture for decision making. If approached intelligently (e.g. state labels represented by a bit-string), logical optimization can be applied through state transition tables and state maps/implication charts. More on this here: https://inst.eecs.berkeley.edu/~cs150/sp00/classnotes/katz-ch9-mod.pdf. to handle idiomatic camera actions. Both of these methods use virtual simulations with user-led events to demonstrate the ability of their tools to derive an idiomatic responses in a short amount of time.
@@
There are more novel approaches to the problem of deriving a shot-list, such as de Lima et al [146] who employ an architecture that simulates four critical filming roles: (1) a Script Writer who observes the context of a current scene and sends information to (2), (2) a Screen-orgrapher who configures the staging of actors and objects for dramatic effect and passes this information to (3) and (4), (3) a Director who extracts important information and uses multiple support vector machines (SVMs) to make decisions on selected shots, and (4) a Cameraman who follows idioms for shooting. This method compounds three 2-D feature matrices into a high dimensional feature space - selecting the most relevant SVM from a pre-trained set to classify the optimal shot position and viewing angle. SVMs work by using kernel maps to remap all features into a high-dimensional matrix. This is done under the assumption that classes of features are not linearly separable until they are compounded into a high-dimension. In practice, normalized positional values represent the environment features, while the perceived emotional state of each actor is represented by the actors’ features, and the actor who is the principal focus is represented within the scene’s features. These are used to for selecting the right SVM to estimate the optimal camera location and angle given the context of the scene. As SVMs may be considered old w.r.t to the current literature, there are likely more suitable approaches for modelling the environment, actor and scene features to classify the optimal camera angle and location from a pre-defined set. For example, with a large enough data set a deep neural networks will generalize better. There are also faster and simpler approaches such as using K-nearest neighbours on the environment, scene and actor features.
@@
Alternatively, CamDroid is a well known state-based camera control architecture, that was introduced by Drucker and Zeltzer [73] as illustrated in Figure 7. This method uses Tool Command Language (Tcl)  [^3] A high-level functional programming language. with TK embeddable language widgets [180] to interface between user-input and controllers, and pre-specified scripts. The pre-specified scripts access an object database (containing information about the local real environment) through application specific processes/object interfaces, and camera modules through camera interfaces; which ultimately determine the camera state and its subsequent action - dependant on a set of filming constraints such as the type of camera motion. Conditional functional frameworks like CamDroid are a classical way of handling automated camera control and are powerful for capturing real-time time-dependant action.
Figure 7: A diagram of the CamDroid architecture, Drucker and Zeltzer [73]

### v3

@@ 2.4 Automated 3-D Capture
Context
Automated 3-D capture is a popular topic in IC research, whether for analysis or for solving inverse graphics problem. Here, the objective is to discover a 3-D representation using 2-D imagery and/or other sensor data. In Figure 4, we illustrate simplified representations of the current approaches to the problem.
@@
(b) Neural Radiance Fields sample points in space to estimate a color and density value for each sample projected along a pixel-ray
(c) Gaussian Splatting uses a point cloud to estimate the position of samples in space and uses this to estimate the color, density scale and rotation of each Gaussian point
@@
A popular capture method is Laser Informed Photogrammetry (e.g., LiDAR) which creates 3-D point-clouds from measurements of point depth relative to a camera’s position. This is now possible on low cost platforms, such as a mobile phone, thanks to the introduction of mobile LiDAR camera sensors. These methods are promising for capturing enclosed spaces, however challenges remain regarding processing information over greater distances as well as limitations associated with noisy data.
Neural Radiance Fields
@@
Recently, neural radiance field (NeRF) methods have been introduced to solve this problem, whereby the aim is to model the rendering processes involved with automated capture. Unlike Photogammtery, which represent scenes as 3-D structures (mesh, voxel-grid, point cloud), NeRF methods represent scene properties as neural representations which are sampled to interpolate views defined by the camera extrinsics in a virtual space. Given a set of images describing a real environment, NeRF networks can reliably learn the visual features of a scene.
@@
NeRFs represent an important breakthrough for cinematographers, and can be adapted to a range of cinematographic tasks. For example, rendering 3-D from 2-D images can avoid costly re-shoots caused by poor lighting, scenery, weather or acting deficiencies [177, 267, 300]. Shots can also be re-worked in post-production, meaning that shot-type, camera jitter, shot duration and focus can also be modified. While the field of NeRF research is relatively new, it has developed rapidly and can now deliver high quality, compact solutions, with the promise of delivering production-ready NeRFs in the near future.
The original NeRF paper [176] demonstrates that, with a good approximation of the camera extrinsic (image position and orientation) and a reliable ray sampler (for sampling points/volumes along a ray) as prerequisites, we can model 3-D scenes that include reflective and transparent materials using a multi-layered perceptron (MLP) neural network. This works by sampling a point along a ray (pertaining to a given pixel in a training image), and using the global 3-D position of that point, $\{x,y,z\}$ and the relative viewing direction of the point, $\{\theta,\phi\}$ as inputs into an MLP that returns an RGB color and an opacity value associated with that point in 3-D space. By accumulating many points along a ray, we can apply a rendering method akin to alpha blending to render each pixel in an image. Through the use of a loss function comparing the rendered image with a ground truth image we can optimize the quality the 3-D scene so that new, unseen views may be generated. Hence, NeRF research falls under the label of novel view synthesis. For readers that are interested in a deeper technical discussion, Appendix A.1 provides a technical introduction to NeRF, how it works and the relevant technical challenges.
Figure 5: Visual comparison of the NeRF and PlenOctrees pipelines used for neural rendering presented by Yu et al [296]. Unlike the original NeRF, NeRF-SH uses spherical harmonics to decode visual features into RGB color values. The PlenOctree model refines a 3-D octree representation to model the geometry of a scene
@@
While this method works, MLPs take long to train and the quality is often poor and reliant on the number of training images/views we have of the 3-D scene. Subsequently, several approaches have been proposed to speed up NeRF computation: multi-resolution hash encodings, voxel grids and voxel-trees [181, 274, 296, 143]. This is particularly relevant for researchers interested in the dynamic NeRF paradigm, as model complexity increases significantly when modeling 4-D space and time. The PlenOctrees representation [296], illustrated in Figure 5, builds off NeRF’s ability to functionally represent a view-dependent scene by, (1) representing the structure of a scene as a sparse-voxel octree [281, 232, 143], and (2) sampling color by involving spherical harmonics. (1) facilitates the rendering step through fast-to-access representations (octrees). Whereas (2) allows us to model view-dependency as octree leaves by mapping the view dependent color values to a sphere’s surface and predicting the coefficients of the spherical harmonic equation (with fixed degree) to return a color value where the pixel-ray intersects the sphere. Subsequently, the rendering speed and learning is optimized significantly, and one could argue novel-view quality has improved too. One caveat is the amount of data a representation will consume (in the magnitude of gigabytes), which perhaps is acceptable for those with required hardware, however there exist more affordable explicit options [47, 223]. Additionally, there is some difficulty with visualizing unbounded scenes as voxel-grids occupy finite space. Yu et al [296] overcomes this by using NeRF++ [303] for rendering out-of-bound scenery. This is accomplished by modeling the foreground and background as separate components.
Gaussian Splatting
@@
With similar motivations, 3-D Gaussian Splatting (GS) has been proposed by Kerbl et al [135] as a means of significantly reducing rendering time by using: (1) a point cloud representation of Gaussian “blobs” with position, covariance, color and opacity properties, which avoids unnecessary computation of empty space present in NeRF models, and (2) tile splatting for faster rendering (NeRF typically renders at $<1$ FPS).
@@
Tile splatting speeds up rendering through the following steps. We highlight the importance of the sorting approach, (3), as this limits NeRFs from achieving similar computational goals.
@@
To train a GS model a scene is initialized with a sparse set of point-clouds that are “densified” during training by cloning, re-scaling and re-positioning blobs to fit the geometry corresponding to a set of training view/image. Blobs that are essentially transparent (low $\alpha$ contribution) are pruned.
Conclusively, GS maybe selected over NeRFs due to their enhanced computational abilities, and in many cases improved performance. However, existing dynamic GS models are much less compact than NeRF alternatives and are comparatively worse at depth estimation [81], due to the differences in depth-based rendering approaches Additionally, as GS is a more recent development than NeRF we are yet to see the same attention on cinematographic based tools, such as scene and camera editing or representational transformations such as GS to mesh. However, the rapid development of this research space means that it is only a matter of time before these limitations are addressed.
@@ 3.1 General Production Applications
In this section we review research relating to camera management and control, visual analysis, cinematographic assistance and workflow optimization.
@@
3.1.1 Computational Language Structures for Scene Analysis, Automated Labeling Schemes and Camera Management
@@
Cinematographic language is used to communicate the state of production and relevant processes. It enables production staff to effectively communicate and react coherently to unforeseen situations. Cinematic theory is often formalized for tasks such as automated camera control and scene analysis. For example, in Jhala and Young [122] state machines are used to model cinematographic shots for camera planning in a virtual environment. In this context, we discuss current approaches automated camera control that introduce or rely on formal cinematic language for inference; we refer to this as idiomatic capture.
@@
A heuristic approach to camera control was introduced by Jhala and Young [121], based on the decompositional partial order causal link planning algorithm, [293]. This formalizes idiomatic capture by linking an approximated scene representation to a predetermined set of camera control responses. A conflict resolution algorithm ranks known actions with conditional operators to determine the optimal action and respective duration. To express idioms, Jhala and Young [121] distinguishes a set of four requirements to mediate contextual differences between different productions: (1) Story Representation, (2) Real World Representation, (3) Rhetorical Coherence and (4) Temporal Consistency. The first two points describe the physical and contextual landscapes of a set, hence it is necessary to have both geometric (or visual) and semantic representations of a production. The third point expresses the need for rhetorical structure to ensure actions are executed decisively, e.g. a hierarchical-model for selecting shot-types. The final point specifies consistency with regards to the temporal aspect of filming.
@@
Similarly, the Declarative Camera Control Language (DCCL) uses a heuristic-based decision tree for idiomatic capture [57], Figure 6. This method sets forth a hierarchical structure for automated shot-composition and camera control by breaking down scenes into idiom-specific frame sequences. Here, the relationship of consecutive frame sequences is dependent on temporal links between idioms. A heuristic evaluator then selects a candidate action by scoring possible responses and evaluating decisions via a decision tree.
@@
Figure 6: Visualizing DCCL for automated camera control: The three stage pipeline uses idioms defined by camera and actor states as inputs to a DCCL compiler. Selected frames are evaluated w.r.t to the “film tree”
@@
Idiomatic language structures have also been used for shot labeling [226] and evaluation [298]. More recently, the Prose Storyboard Language (PSL) [226] has been proposed as a high-level language for describing shots relative to visual composition and idiomatic camera control. This is demonstrated in Alfred Hitchcock’s North by Northwest, where a prompt not only describes the shot-type and foreground, mid and background compositions (separately) but the type of transition as well. As PSL uses AND/OR decision trees it can be easily configured to add/modify labels. While the authors have not tested this on an automated labeling scheme, there is cause for such investigation as idiom-based languages might overlook newly adopted idioms and foreign capture practices.
@@
Embedded Constrained Patterns (ECP) is a dedicated query language, conscious of physical cinematographic limitations [285]. Unlike PSL, ECP is more comprehensive in its description - labeling montages rather than individual shots. Labels are assigned using the following descriptors:
@@
Wu and Christie [285] propose ECP alongside a automated search algorithm for optimal shot-sequence description. The search method is partitioned into two stages, (1) build a cache of valid solutions, and (2) apply a double recursive depth search to choose optimal description from the set of valid solutions. During stage (1) valid solutions are separated into three sets: (i) frames that satisfy constraints pre-defined by ECP, (ii) frame couples that satisfy the ECP’s relational constraint, and (iii) frame sequences. In the second stage, the search process iterates over all frames in each sequence where subsequent frames are validated as part of the local or global sequence. Alongside the technical implementation of ECP, Wu and Christie [285] suggested plans for integration with Unity 3-D to apply ECP to their montages; which we see as a positive contribution towards the production pipeline.
Additionally, languages constructed for automated camera control are not easily comparable. The choice of language structure can vary drastically depending on the application. While there is no current standard for general cinematography, we observe interest from Movie Labs, who look to define a set of computational linguistic structures as a standard for cloud-based workflow optimization [142]. The shift to cloud-based computing is indicative of another paradigm of production. Notably, as assets and tools are usually stored or executed offline changes to shared resources can be difficult to monitor on one platform. This could be facilitated by file management approaches which mimic, for example, code-sharing practices. However, this may be difficult to scale for productions requiring a large number of assets or productions which involve external international collaborators who may be un-knowledgeable of standard practices.
@@ 3.1.2 Directive Assistants
Converting most cinematographic concepts into controllers presents a challenging task [58]. Notions of shot compositions, shot types and shot transitions are bound by real world problems such as the cost of production, physics and the topology of a set. Thus, AI directive assistants (DAs) have been introduced to alleviate these production challenges. Current IC research addresses tasks such as deriving shot lists, shot plans, optimizing camera placement and controlling robotic camera rigs.
@@
There is a tendency to use semantic representations of a work environment to provoke DAs. For example, through DCCL, Christianson et al [57] demonstrates a heuristic approach that relies on idiomatic-based practice for decision making on shot composition. Similarly, He et al [100] discuss a different heuristic approach, utilizing a set of finite state machines [^1] A state-based control architecture for decision making. If approached intelligently (e.g., state labels represented by a bit-string), logical optimization can be applied through state transition tables and state maps/implication charts. More on this here: https://inst.eecs.berkeley.edu/~cs150/sp00/classnotes/katz-ch9-mod.pdf. to handle idiomatic camera actions. Both of these methods use virtual simulations with user-led events to demonstrate the ability of their tools to derive an idiomatic responses in a short amount of time.
@@
There are more novel approaches to the problem of deriving a shot-list, such as de Lima et al [156] who employ an architecture that simulates four critical filming roles: (1) a Script Writer who observes the context of a current scene and sends information to (2), (2) a Screen-orgrapher who configures the staging of actors and objects for dramatic effect and passes this information to (3) and (4), (3) a Director who extracts important information and uses multiple support vector machines (SVMs) to make decisions on selected shots, and (4) a Cameraman who follows idioms for shooting. This method compounds three 2-D feature matrices into a high dimensional feature space - selecting the most relevant SVM from a pre-trained set to classify the optimal shot position and viewing angle. SVMs work by using kernel maps to remap all features into a high-dimensional matrix. This is done under the assumption that classes of features are not linearly separable until they are compounded into a high-dimension. In practice, normalized positional values represent the environment features, while the perceived emotional state of each actor is represented by the actors’ features, and the actor who is the principal focus is represented within the scene’s features. These are used to for selecting the right SVM to estimate the optimal camera location and angle given the context of the scene. As SVMs may be considered old w.r.t to the current literature, there are likely more suitable approaches for modeling the environment, actor and scene features to classify the optimal camera angle and location from a pre-defined set. For example, with a large enough dataset a deep neural networks will generalize better. There are also faster and simpler approaches such as using K-nearest neighbours on the environment, scene and actor features.
@@
Alternatively, CamDroid is a well known state-based camera control architecture, that was introduced by Drucker and Zeltzer [76] as illustrated in Figure 7. This method uses information on the local real environment to inform a sequence of actions to be taken for optimal capture. Whereby, different visual queues or user prompts will trigger a sequence of camera actions. Conditional functional frameworks like CamDroid are a classical way of handling automated camera control and are powerful for capturing real-time time-dependent action.
Figure 7: Visualizing CamDroid for camera control: A state-based architecture that relies on user inputs and environment constraints to modify camera parameters for an action sequence